# Computing Certificates of Strictly Positive Polynomials in Archimedean Quadratic Modules

### v1

@@ 6.1.1 Certificates of strictly positive polynomials
The objective of this benchmark is to compare the quality of certificates and the time required by RealCertify  [^4] We use the command multivsos_interval(f, glist=generators, gmp=true). and our implementation. This benchmark uses as generators the polynomial set of the form $\{p,-p\}$ where $p:=\sum_{i=1}^{2n}{(x-a_{i})}$ and computes certificates of strictly positive polynomials of the following kinds: a linear polynomial of the form $x-a$ (left_strict in Table 1), a linear polynomial of the form $-(x-a)$ (right_strict), the product of the generators $-p^{2}+35$ (lifted_prods), and the Archimedean polynomial added to the generators for RealCertify plus $10$ (arch_poly).
We produced 100 tests for each of these kinds of problems, and Table 1 shows the results obtained. Each problem contains two rows: the first row uses RealCertify adding an Archimedean polynomial to the set of generators [^5] The RealCertify package uses the algorithms in [9] which require an explicit assumption about the Archimedeaness of the quadratic module; for this reason we extend the set of generators with an Archimedean polynomial as otherwise RealCertify does not find certificates.; the second describes the results obtained by our implementation. To compare the quality of the certificates obtained, we gather the degrees of the sum-of-squares multipliers obtained. We report these in the second column “Max. Poly Degrees” collecting the maximum value of all the tests, the average and median. The third column “Time” reports the time in seconds for the average, median, and total of the 100 tests. Finally, the fourth column reports the number among 100 tests, for which certificates were computed.
Problem Max Poly. Degrees Time (seconds) Successful
@@
left_strict 8 5.8 6 0.041 0.036 3.557 87/100
14 7.16 6 0.092 0.057 9.185 100/100
right_strict 8 5.775 6 0.041 0.036 3.287 81/100
8 6.88 8 0.064 0.058 6.411 100/100
lifted_prods 8 5.775 6 0.007 0.007 0.548 74/100
8 6.88 6 0.064 0.057 6.387 100/100
arch_poly 8 5.932 6 0.039 0.037 1.667 43/100
8 7.06 6 0.085 0.07 8.493 100/100
@@
As the table shows, even though our implementation runs slower than RealCertify, it is able to find certificates for all the examples in the benchmark; in contrast, RealCertify cannot compute certificates for some of these examples. Additionally, our implementation can compute certificates using only the elements of the original set of generators given, whereas RealCertify requires a polynomial to witness the Archimedeaness of the quadratic module. The Archimedean polynomial is always a member of compact quadratic modules in the univariate case, the resulting certificates would include an additional element in the set of generators for which certificates need to be computed explicitly to obtain certificates in the original generators.
6.1.2 Strictly positive polynomials close to the semialgebraic set
For the second benchmark, we use a parameterized monogenic generator $\{(1-x^{2})^{k}\}$. It is well known in the literature [20, 13] that the degree of the certificates depends on the degree of the generators, as well as the separation of the zeros of the input polynomial and the endpoints of the semialgebraic set associated with the set of generators. We show with our experimental results that our approach and heuristics can effectively handle this problem.
The benchmark evaluates the degree obtained by our implementation and compares it with RealCertify. Since the algorithm discussed in Section 5 to compute certificates of strictly positive polynomials can handle any set of generators, we use the monogenic generator $\{(1-x^{2})^{k}\}$ where $k$ is an odd natural number. The semialgebraic set for this generator is the interval $[-1,1]$. We use different values of $\epsilon>0$ and $k$ of the strictly polynomial $(1+\epsilon)+x$ over $\operatorname{S}(\{(1-x^{2})^{k}\})$ to report the degrees of the certificates obtained and the time (in second) taken in Table 2 below.
$\epsilon$ $k$ RealCertify Ours
Time Degree Time Degree
$\frac{1}{2}$ 13 0.09 26 0.072 26
15 0.199 30 0.075 30
17 0.254 34 0.077 34
19 0.288 38 0.078 38
21 0.382 42 0.08s 42
$\frac{1}{3}$ 13 0.175 26 0.05 26
15 0.205 30 0.052 30
17 0.247 34 0.056 34
19 0.297 38 0.059 38
21 0.391 42 0.074 42
@@ 6.1.2 Strictly positive polynomials close to the semialgebraic set
6.2 Certificates of globally strictly positive polynomials
This benchmark is used to test our implementation of the extended Lasserre’s method discussed in Section 4. It uses two strictly positive polynomials by adding a positive constant to the Motkzin and Robinson polynomials. These polynomials are, respectively, of the form $x^{6}+y^{4}z^{2}+y^{2}z^{4}-3x^{2}y^{2}z^{2}+1/d$ and $x^{4}y^{2}+y^{4}z^{2}+x^{2}z^{4}-3x^{2}y^{2}z^{2}+1/d$ where $d$ is a positive constant.
Figure 1: Performances of implementation of Algorithm 3 for modified Motkzin and Robinson polynomials
Figure 1 reports the parameter $m$ and the time (in second) used by the implementation. The parameter $m$ is the minimal degree for $f_{\epsilon}$ in (3) to become a sum of squares. The reader should note that as $d$ (horizontal axis) increases, the values of $m$ and time also increase.
@@ 7 Future work
We conclude this paper by listing several potential research problems we want to further investigate: (1) degree estimation for the certificates constructed by our method, and this may lead to interesting results for effective Putinar’s Positivstellensatz; (2) extending our method to deal with polynomials which are non-negative over $\operatorname{S}(G)$; (3) complexity analysis of Algorithm 4 and further optimization on our implementation for it; (4) identify sub cases and conditions to obtain simpler certificates.
@@ 8 Acknowlegements
The research of Weifeng Shang and Chenqi Mou was supported in part by National Natural Science Foundation of China (NSFC 12471477) and that of Jose Abel Castellanos Joo and Deepak Kapur was supported in part by U.S. National Science Foundation (CCF-1908804).

### v2

@@ 7.1.1 Certificates of strictly positive polynomials
The objective of this benchmark is to compare the quality of certificates and the time required by RealCertify  [^5] We use the command multivsos_interval(f, glist=generators, gmp=true). and our implementation. This benchmark uses as generators the polynomial set of the form $\{p,-p\}$ where $p:=\sum_{i=1}^{2n}{(x-a_{i})}$ and computes certificates of strictly positive polynomials of the following kinds: a linear polynomial of the form xax-a (left_strict in Table 1), a linear polynomial of the form (xa)-(x-a) (right_strict), the product of the generators $p^{2}+35$ (lifted_prods), and the Archimedean polynomial added to the generators for RealCertify plus $10$ (arch_poly).
We produced 100 tests for each of these kinds of problems, and Table 1 shows the results obtained. Each problem contains two rows: the first row uses RealCertify adding an Archimedean polynomial to the set of generators [^6] The RealCertify package uses the algorithms in Magron and Safey El Din (2018a) which require an explicit assumption about the Archimedeaness of the quadratic module; for this reason we extend the set of generators with an Archimedean polynomial as otherwise RealCertify does not find certificates.; the second describes the results obtained by our implementation using the original generators for each benchmark problem. To compare the quality of the certificates obtained, we gather the degrees of the sum-of-squares multipliers obtained. We report these in the second column “Max Poly Degrees” collecting the maximum value of all the tests, the average and median. The fourth column “Time” reports the time in seconds for the average, median, and total of the 100 tests. Finally, the last column reports the number among 100 tests, for which certificates were computed.
Problem Implementation Max Poly Degrees Time (seconds) Successful
@@
left_strict RealCertify 6 5.573 6 0.039 0.037 3.458 89/100
Ours 8 6.62 7 0.05 0.05 5.045 100/100
right_strict RealCertify 8 5.5 6 0.045 0.04 3.597 80/100
Ours 10 6.84 7 0.065 0.058 6.505 100/100
lifted_prods RealCertify 4 4.0 4 0.014 0.01 1.365 100/100
Ours 8 5.12 5 0.162 0.03 16.168 100/100
arch_poly RealCertify 8 5.236 6 0.045 0.039 2.459 55/100
Ours 10 6.74 7 0.088 0.073 8.772 100/100
@@
As the table shows, even though our implementation runs slower than RealCertify, it is able to find certificates for all the examples in the benchmark; in contrast, RealCertify cannot compute certificates for some of these examples. Additionally, our implementation can compute certificates using only the elements of the original set of generators given, whereas RealCertify requires a polynomial to witness the Archimedeaness of the quadratic module. The Archimedean polynomial is always a member of compact quadratic modules in the univariate case, and the resulting certificates by RealCertify would include an additional element in the set of generators for which certificates need to be computed explicitly if one wants certificates in the original generators.
7.1.2 Comparison of degree of certificates
For the second benchmark, we use a parameterized monogenic generator $(1-x^{2})^{k}$ where $k$ is an odd number. It is well known in the literature Stengle (1996); Powers and Reznick (2000) that the degree of the certificates depends on the degree of the generators, as well as the separation of the zeros of the input polynomial and the endpoints of the semialgebraic set associated with the set of generators. We show with our experimental results with this kind of benchmark that our approach and heuristics can effectively handle this problem.
The benchmark evaluates the degree obtained by our implementation and compares it with RealCertify. The semialgebraic set for this generator is the interval [1,1][-1,1]. We use different values of $\epsilon>0$ and $k$ of the strictly positive polynomial $(1+\epsilon)+x$ over [1,1][-1,1] to report the degrees of the certificates obtained in $\operatorname{QM}((1-x^{2})^{k})$ and the time (in seconds) taken in Table 2 below. Additionally, we provide complexity estimates for this benchmark from Stengle (1996). We notice both implementations produce certificates within the theoretical bounds.
$\epsilon$ $k$ RealCertify Ours Theoretical bounds
Degree Time Degree Time Lowerbound Upperbound
$\frac{1}{2}$ 13 26 0.094 26 0.02 2 94
15 30 0.215 30 0.02 2 106
17 34 0.247 34 0.021 2 118
19 38 0.312 38 0.022 2 130
21 42 0.361 42 0.023 2 142
$\frac{1}{3}$ 13 26 0.181 26 0.02 6 128
15 30 0.182 30 0.02 6 146
17 34 0.252 34 0.022 6 162
19 38 0.319 38 0.024 6 180
21 42 0.389 42 0.026 6 196
@@ 7.1.2 Comparison of degree of certificates
7.2 Certificates of globally strictly positive multivariate polynomials
This benchmark is used to test our implementation of the extended Lasse-rre’s method discussed in Section 4. The quadratic module in consideration is generated by the polynomial $g:=10-x^{2}-y^{4}-z^{8}$. The quadratic module $\operatorname{QM}(g)$ is Archimedean as it contains the Archimedean polynomial $14-x^{2}-y^{2}-z^{2}$; its certificates are
$\begin{split}&14-x^{2}-y^{2}-z^{2}=\frac{3854x^{2}}{44027}+\frac{3854y^{4}}{44027}+y^{2}+\left(1-y^{2}\right)^{2}+\frac{8047692458107z^{8}}{1282595968637408}\\ &+\frac{273z^{6}}{26240}+2z^{4}+\frac{123}{160}\left(z-\frac{55z^{3}}{82}\right)^{2}+\frac{910375997}{5070233600}\left(z^{2}-\frac{591644335z^{4}}{910375997}\right)^{2}\\ &+\frac{49514}{44027}\left(\frac{220135z^{4}}{3168896}-\frac{12459641z^{2}}{15844480}+1\right)^{2}+\left(1-z^{4}\right)^{2}+\left(\frac{3854}{44027}+1\right)\cdot g.\end{split}$
For the input polynomials, we use two strictly positive polynomials by adding a positive constant to the Motkzin and Robinson polynomials. These polynomials are, respectively, of the form $x^{6}+y^{4}z^{2}+y^{2}z^{4}-3x^{2}y^{2}z^{2}+\frac{1}{d}$ and $x^{4}y^{2}+y^{4}z^{2}+x^{2}z^{4}-3x^{2}y^{2}z^{2}+\frac{1}{d}$ where $d$ is a positive constant. We encoded the problem in RealCertify and used several values for its configuration parameter relaxorder; however, the package cannot find certificates for this benchmark.
Figure 1: Performances of implementation of Algorithm 3 for lifted Motkzin and Robinson polynomials by $\frac{1}{d}$.
Figure 1 reports the parameter $r$ and the time (in seconds) used by the implementation. The parameter $r$ is the minimal degree for $f_{\epsilon}$ in (6) to become a sum of squares. The reader should note that as $d$ (horizontal axis) increases, the values of $r$ and time also increase.
@@ 8 Future work
We conclude this paper by listing several potential research problems to investigate: (1) degree estimation for the certificates constructed by our method; a consequence of this line of research could be new interesting results for effective Putinar’s Positivstellensatz; (2) extend the proposed method to deal with polynomials which are non-negative over $\operatorname{S}(G)$; (3) complexity analysis of Algorithm 4 possibly leading to optimization of our implementation; (4) identify sub cases and conditions to obtain lower degree certificates.
@@ 9 Acknowlegements
The research of Weifeng Shang and Chenqi Mou was supported in part by National Natural Science Foundation of China (NSFC 12471477). Jose Abel Castellanos Joo and Deepak Kapur were supported in part by U.S. National Science Foundation awards (CCF-1908804) and (CCF-2513374).